The Express Banquet has two restaurants that are open 24 hours a day. Fixed costs for the two restaurants together total $509,000 per year. Service varies from a cup of coffee to full meals. The average sales bill per customer is $9.50. The average cost of food and other variable costs for each customer is $3.80. The income tax rate is 30​%. Target net income is $122,500.

Required

1. Compute the revenues needed to earn the target net income.

2. How many customers are needed to earn net income of $122,500​? How many customers are needed to break​ even?

3. Compute net income if the number of customers is 170,000.

Requirement 1. Using the equation​ method, select the basic formula used to compute the revenues needed to earn the target net income.
Answer #1
1.) Sales Revenue $ 1,140,000 =(509000+(122500/70%))/60%
( Fixed cost + Pretax profit )/ Contribution margin ratio
Working:
Contribution margin ratio 60% =(9.5-3.8)/9.5
2.) Customers needed    120,000 =1140000/9.5
( Taget sales Revenue / average sale bill)
Breakeven point-Customers        89,298 =509000/(9.5-3.8)
( Fixed cost / Contribution per customer )
3.) Amount $
Sales Revenue ( 170,000 x 9.50 ) 1,615,000
Less: Variable costs ( 170,000 x 3.80 )    646,000
Contribution margin    969,000
Less: Fixed costs    509,000
Earnings before tax    460,000
Less: Income tax @30%    138,000
Net Income    322,000